OpenStudy (anonymous):

5/(2x+6) - 2= (1-8x)/4x ?

OpenStudy (anonymous):

First take the LCM on both sides and solve for x you will get x= 1/3

hero (hero):

Multiply -2 by 4x/4x

hero (hero):

\(\frac{5}{2x+6} - \frac{8x}{4x} = \frac{1-8x}{4x}\)

hero (hero):

Add 8x/4x to both sides: \(\frac{5}{2x+6} = \frac{1-8x}{4x} + \frac{8x}{4x}\)

hero (hero):

\(\frac{5}{2x+6} = \frac{1}{4x}\)

hero (hero):

20x = 2x + 6 20x - 2x = 6 18x = 6 x = \(\frac{6}{18}\) x = \(\frac{1}{3}\)